You bought a convertible bond issued by HiTech Corp which has a conversion ratio of 50 common shares for each $1,000 bond. At what stock price per share would you make a profit (“in the money”) if you bought the bond at par? What would you expect the bond to sell for in the market if the stock trades at $28 per share?
Answer #1

a: Conversion ratio = Par value/Share price

50= 1000/Share price

Share price = 1000/50 = $20

b: Price of bond = Conversion Value = Conversion Ratio × Current Share Price

= 50*28 = $1400